Cintas Corporation has debt with both a face and a market value of $1,020,000. This debt has a coupon rate of 6 percent and pays interest annually. The expected earnings before interest and taxes are $820,000, the tax rate is 25 percent, and the unlevered cost of capital is 10.04 percent. What is the firm's cost of equity?
Answer #1

Step 1: VU = EBIT * (1-Tax rate) / Cost of capital

VU = 820000 (1-.25) / .1004

VU = $6125498

Step 2: VL = VU + (Tax rate * Debt)

VL = 6125498 + (.25*1020000)

VL = $6380498

Step 3: VE = VL-VD

VE = 5360498

Step 4: Re = RU + (RU-RD) * (D/E) * (1-Tax rate)

RU = Cost of capital, RD = Cost of debt

Re = .1004 + (.1004-.06) * (1020000/5360498) * (1-.25)

Re = .1004 + .0057655

Re = .1061655

Cost of Equity (Re) = 10.62%